Use the two-way table below to answer the following question. Kids Total 10 15 Cars 2 TotalS 30 Question 1P< 3 cars | 2 to 4 kids) Question 2 P( 3 kids or2 cars) Click to Grade Your Work
Answer #1

1)P(<3 cars given 2 to 4 kids)=P(<3 cars and 2 to 4 kids)/P(2 to 4 kids)

=(1+2+2+2+1+3)/(5+5+5)=11/15

2)

P(>3 kids or >=2 cars)=1-P(<=3 kids and < 2 cars)=1-(4/30)=26/30=13/15